I am helping in after school tutoring and I want to be sure I am correct-
when you do an equation such as : 5z squared + 9z squared = 504
what do you do with the exponents?

Answer by wuwei96815(245)

In addition the exponents do not change.

5z^2 + 9z^2 = 504
14z^2 =504
z^2 = 504/14 = 36
z = 6